OpenStudy (tainted):

Explain, in complete sentences, how you would divide the following expression:^11√x^5/8√x^3

OpenStudy (abb0t):

Can you re-write it. I don't quite understand what ^11 means. Or is it x^5/11??

OpenStudy (tainted):

here is the problem

OpenStudy (abb0t):

re-write the roots in fractions. So i think if i remember my root rules correctly the numerator is \[x ^{\frac{ 11 }{ 5 }}\] and the denom is \[x ^{\frac{ 8 }{ 3 }}\] now you can subtract to remove like terms. Kinda like how \[\frac{ x^{2} }{ x }=x\] you know that because 2-1 = 1 so you just have the 1 power. Same thing applies. \[\frac{ 11 }{ 5 }-\frac{ 8 }{ 3 }\] and you should be able to solve for you're x now.

OpenStudy (abb0t):

\[answer: x ^{\frac{ 11 }{ 5 }-\frac{ 8 }{ 3 }}\]